DIRC Operations Weekly Report
Period: Wednesday 2007/8/01 - Wednesday 2007/8/08
Oncall: Justine Serrano
Fast Monitoring: Rolf Andreassen, Justine Serrano
Luminosities this week
PEP peak lumi: ~1.05 x 1034
(last week: ~1.02 x 1034)
PEP delivery: 3.66 fb-1
(last week: 4.64 fb-1)
BaBar integrated lumi:3.21 fb-1
(last week: 4.5 fb-1)
Run 6 total delivery: 72.45 fb-1
Run 6 total BaBar recorded: 69.66 fb-1 total, including 4.18 fb-1 off-peak data
PEP activities/issues this week
Aug 1: Install helium lines to look for an IR2 vacuum leak.
Aug 2: No beam during most of the day to fix a water leak in the Linac.
Aug 3: starting to have problem with the RF station 4-3.
BaBar activities/issues this week
DIRC activities/issues this week
Aug 2: One christmas tree showed up in sector 5 during cosmic data taking. We turn off channel 6 in sector 5 (Set V0 and V1 at 110V) to avoid trip. There was another in sector 11, but it disappeared after turn off and on the high voltage (channels 5 and 10).
Aug 2: Masked out a noisy channel in sector 10.
Aug 6: Neutron counters' power supply tripped again, Matt replaced the power supply.
See DIRC logbook for details
Lumi and Background this week
- Next wednesday, UPS battery replacement will require Matt to disable SOB dump valve.
- Replace four DFBs with goups of 8 dead or flaky channels:
- sector 0, DFB 0.
- 2 in sector 1, DFB 1 and DFB 13.
- sector 3, DFB 6.
- sector 8, DFB 0.
- Replace air filters on Wiener crates.
- Investigate the crate of sector 8.
- Unplug Christmas tree in Sector 5.
- Investigate Christmas tree in Sector 11.
Justine, Jose and Rolf rotating backup & on-call commissioning duties.